# Invertible Consistency Distillation for Text-Guided Image Editing in Around 7 Steps

### v1

@@ 3.4 Dynamic Classifier-Free Guidance Facilitates Inversion
As previously discussed, dynamic guidance [31, 32] provides promising results for both faithful and diverse text-to-image generation. In this work, we reveal that dynamic CFG is also an effective technique for improving inversion accuracy as shown in Figure 4a. Below, we delve into the questions when and why dynamic guidance might facilitate image inversion while preserving the generative performance. To answer these questions, we conduct experiments using Stable Diffusion 1.5 with DDIM solver for $50$ steps and maximum CFG scale set to $8.0$.
@@ 4.2 Text-guided image editing
Figure 8: Visual examples of different editings provided by our and baseline approaches using SD1.5.
Figure 9: Quantitative comparisons between different approaches based on the SD1.5 model: automatic metrics (left) and human evaluation (right).
@@
Table 3: Automatic metrics (upper) and human evaluation (bottom) for iCD-XL and ReNoise [9].
@@ Appendix A Training details of iCD
First, following [6], we preliminary distill the CFG using a conditional embedding added to the time step embedding. The goal is to use different $w$ and apply dynamic techniques during inference after consistency distillation. We perform CFG distillation for the following guidance scales: $1,8,12,16,20$ for the SD1.5 and $1,4,6,9,10,12,13,16,18,20$ for the SDXL. At this stage, the model successfully approximates the guided teacher without hurting its performance.
Then, we perform multi-boundary consistency distillation using LoRA adapters with a rank of $64$ following [39]. We train forward and reverse models in parallel starting from the same teacher: CFG distilled SD1.5 or SDXL. For the SD1.5 model, we use a global batch size of $512$, and for the SDXL $128$. All models converge relatively fast, requiring about $6$K iterations with a learning rate $68e{-}$ 6.
We find that $4$ and $3$ steps of the forward and reverse models provide promising generation and inversion quality. The regularization coefficients for the forward and reverse losses are equal to $\lambda_{f}{=}1.5$ and $\lambda_{r}{=}1.5$, respectively.
The iCD-SD1.5 models are trained for ${\sim}36$h and the iCD-XL ones for ${\sim}72$h on $8$ NVIDIA A100 GPUs. We present the training dynamics for iCD-SD1.5 in Figure 11.
@@
  • $3$ steps $\tau=0.7$: reverse model $\left[339,679,999\right]$; forward model $\left[19,339,679\right]$;
@@ Appendix B Image generation with iCD
Table 4: Text-to-image performance of the SD1.5 model in terms of FID, CLIP score and ImageReward using $w=8$ using $5$K prompts from the MS-COCO dataset.
@@
Figure 13: The human evaluation interface of the text-guided image editing problem.
@@ Appendix C Image inversion
Table 5: More image inversion results with iCD-SD1.5 using $3$ and $4$ steps. d.CFG employs $\tau=0.7$.
@@ Appendix D Text-guided image editing
We present the instruction for the assessors in Figure 13. We calculate the confidence interval using bootstrap methods, splitting the human votes into $1,000$ subsets and then averaging the results and calculating the standard deviation.
Figure 12 demonstrates the comparison between two configurations of the iCD ($3$ and $4$ steps). We observe that both configurations perform similarly, with a slight preference of the $4$-step configuration. We present additional visual results on image editing using the iCD-SD1.5 model in Figure 18 and the iCD-XL model in Figure 19.

### v2

@@ 3.4 Dynamic Classifier-Free Guidance Facilitates Inversion
As previously discussed, dynamic guidance [44, 45] provides promising results for both faithful and diverse text-to-image generation. In this work, we reveal that dynamic CFG is also an effective technique for improving inversion accuracy as shown in Figure 4b. Below, we delve into the questions when and why dynamic guidance might facilitate image inversion while preserving the generative performance. To answer these questions, we conduct experiments using Stable Diffusion 1.5 with DDIM solver for $50$ steps and maximum CFG scale set to $8.0$.
@@ 4.2 Text-guided image editing
Figure 8: Image editing examples produced by our method (iCD-SD1.5) and the baseline approaches.
Figure 9: Quantitative comparisons between different editing approaches based on SD1.5: automatic metrics (left) and human preference study (right).
@@
Table 3: Automatic metrics (top) and human evaluation (bottom) for iCD-XL and ReNoise [24].
@@ Appendix A Training details of iCD
First, following [19], we preliminary distill classifier-free guidance using a conditional embedding added to the time step embedding. The goal is to use different $w$ and apply dynamic techniques during inference after consistency distillation. We perform CFG distillation for the following guidance scales: $1,8,12,16,20$ for SD1.5 and $1,4,6,9,10,12,13,16,18,20$ for SDXL. At this stage, the model successfully approximates the guided teacher without hurting its performance.
Then, we perform multi-boundary consistency distillation using LoRA adapters with a rank of $64$ following [51]. We train forward and reverse adapters in parallel starting from the same teacher: CFG distilled SD1.5 or SDXL. For the SD1.5 model, we use a global batch size of $512$, and for the SDXL $128$. All models converge relatively fast, requiring about $6$K iterations with a learning rate $68e{-}$ 6.
We find that the forward and reverse models provide promising generation and inversion quality for $3$ or $4$ steps. The regularization coefficients for the forward and reverse preservation losses are $\lambda_{f}{=}1.5$ and $\lambda_{r}{=}1.5$, respectively.
The iCD-SD1.5 models are trained for ${\sim}36$h and the iCD-XL ones for ${\sim}68$h on $8$ NVIDIA A100 GPUs. We present the training dynamics in terms of FID and reconstruction MSE for iCD-SD1.5 in Figure 11.
@@
  • $3$ steps $\tau=0.7$: reverse model $\left[339,699,999\right]$; forward model $\left[19,339,699\right]$;
@@ Appendix B Image generation with iCD
Table 4: Text-to-image performance of the SD1.5 model in terms of FID-$5$K, CLIP score and ImageReward for $w=8$ using $5$K prompts from the MS-COCO dataset.
@@
Figure 13: The human evaluation interface for the text-guided image editing problem.
@@ Appendix C Image inversion
Table 5: More image inversion results for 3- and 4-step iCD-SD1.5. Dynamic CFG uses $\tau=0.7$.
@@ Appendix D Text-guided image editing
We present the annotation interface for the assessors in Figure 13. We calculate the confidence interval using bootstrap methods, splitting the human votes into $1,000$ subsets and then averaging the results and calculating the standard deviation.
Figure 12 compares two iCD configurations ($3$ and $4$ steps). We observe that both configurations perform similarly, with a slight preference for the $4$-step configuration. We present additional visual results on image editing using the iCD-SD1.5 model in Figure 18 and the iCD-XL model in Figure 19.